ACCOMPANYING FILE FOR THE DISTRIBUTION OF Scilab

Scilab (hereinafter referred to as SCILAB) has been designed and produced by the “Scilab group” mainly composed by researchers from the META2 research project at INRIA “Institut National de Recherche en Informatique et Automatique - Domaine de Voluceau - Rocquencourt - B.P. 105 - 78153 - LE CHESNAY Cedex - FRANCE”.

INRIA holds all ownership rights relating to SCILAB.

SCILAB is registered at “Agence pour la protection des programmes (APP)” under the number 93-27-011-00.

1 Preamble

SCILAB is currently being developed and INRIA desires that it can be used by the scientific community in a spirit of friendship and cooperation. To this end, INRIA has decided to have SCILAB distributed in source code format by FTP. However, copyright subsists on the part of SCILAB which has been developed at INRIA. SCILAB’s numeric library makes use of several public domain or freely - distributed (and copyrighted by other organizations) programs.

2 Extent of the rights granted by INRIA to the users of SCILAB

INRIA freely grants the right to use, modify and integrate SCILAB in another program provided that :

-this copyright notice appears in all copies

-acknowledgement of receiving SCILAB is made by returning the agreement below Any modification to SCILAB must be brought to the notice of any other user of the modified SCILAB.

3 Reproduction of SCILAB

INRIA grants any user of SCILAB the right to reproduce it in order to circulate it in accordance with the same purposes and conditions as those defined above.

Any copy of SCILAB and/or relevant documentation must include reference to the ownership of INRIA and the present file.

The user undertakes not to carry out any paying distribution of SCILAB.

4 Agreement

INRIA grants to user name : address :

a license to use, copy and modify SCILAB under the terms expressed above. SCILAB is provided ‘as is’, without express or implied warranty.

INRIA will not be responsible for the correction of any bugs or other deficiencies . In no event INRIA shall be liable for any damage of any kind, including special, indirect or consequent damages, arising in connection with the use of SCILAB.

Return address :

Scilab/Meta2 - INRIA Bat. 12

Rocquencourt - B.P. 105

78153 LE CHESNAY Cedex

FRANCE

5 Email address

email : Scilab@inria.fr

Date: Signature: